TFTOPL(1) TFTOPL(1) NAME tftopl−convertTeXfontmetric(tfm)filestopropertylists SYNOPSIS tftopl[−charcode-format=format][−verbose]tfmname[.tfm][plname[.pl]] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof TEXcanbefoundintheinfofileormanualWeb2C:ATeXimplementation. Thetftoplprogramtranslatesa(program-oriented)fontmetricfiletoa(human-oriented)prop-ertylistfile.Thus,youcaneditthecontentsoftheTFMfiles,ifthefontdesignerhasnotdone hisorherjobproperly,orifyou’reencounteringstrangedifficulties,orifyou’rejustcurious.It alsoservesasaTFM-filevalidatingprogram,i.e.,ifnoerrormessagesaregiven,theinputfileis correct. Theplnameisextendedwith.plifitlacksasuffix.Ifplnameisnotgiven,standardoutputis used.Likewise,tfmnameisextendedwith.tfmifitlacksasuffix.Also,pathsearchingisdone forthe.tfmfileusingtheenvironmentvariableTEXFONTS.Ifthatisnotset,tftoplusesthe systemdefault. Seetex(1)forthedetailsofthesearching. OPTIONS Theargumentformatto−charcode-formatspecifieshowcharactercodesareoutputinthePL file.Bydefault,onlylettersanddigitsareoutputusingtheCintegercode(i.e.,inASCII);the othersareoutputinoctal.(Unlessthefont’scodingschemestartswithTeXmathsyorTeX mathex,inwhichcaseallcharactercodesareoutputinoctal.)Ifformatisascii,allcharacter codesthatcorrespondtographiccharacters,besidestheleftandrightparentheses,areoutputin ASCII.Finally,ifformatisoctal,allcharactercodesareoutputinoctal. Withoutthe−verboseoption,tftoploperatessilently.Withit,abannerandprogressreport areprintedonstdout. SEEALSO pltotf(1), DonaldE.Knuth,TEXware. DonaldE.Knuth,TheMETAFONTbook(VolumeCofComputersandTypesetting),Addison-Wes-ley,1986,ISBN0-201-13445-4. AUTHORS DonaldE.Knuthwrotetheprogram.ItwaspublishedaspartoftheTEXwaretechnicalreport, availablefromtheTEXUsersGroup.HowardTrickeyandPavelCurtisoriginallyporteditto Unix. Web2C7.32x 17December1994 1